OpenStudy (anonymous):

8. Solve the following trinomial: + 9x + 14 = 0

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

\[x^2+7x+2x+14=0\] \[x(x+7)+2(x+7) = 0 \] \[(x+7)(x+2)=0\] either x+7=0 or x+2=0 \[x+7=0,x=-7\] \[x+2=0,x=-2\]

OpenStudy (anonymous):

(x+2)(x+7) x= -2 and x= -7